OpenStudy (anonymous):

Is (2,4) a solution to the equation y = x – 2?

OpenStudy (lgbasallote):

substitute x = 2 and y = 4 into your equation

OpenStudy (lgbasallote):

what do you get?

OpenStudy (anonymous):

im honestly not the smartest person in math so i have no freaking idea

OpenStudy (anonymous):

ive tried to solve but but im lost

OpenStudy (lgbasallote):

substitute 2 into x so y = 2 - 2 then substitute 4 into y so 4 = 2 - 2 do you get the substitution?

OpenStudy (anonymous):

no but i wish i did.

OpenStudy (lgbasallote):

y = x - 2 x = 2 and y = 4 so you REPLACE x with 2 and REPLACE y with 4

OpenStudy (anonymous):

oh okay but you didnt have to go all caps lock on me now. jeez just asking for help thats all.

OpenStudy (lgbasallote):

i was capsing to stress the words lol

OpenStudy (lgbasallote):

people notice capitalized words more

OpenStudy (anonymous):

Oh well that does make sense lol

OpenStudy (lgbasallote):

anyway 4 = 2 - 2 is that true? is 4 equal to 2 minus 2?

OpenStudy (anonymous):

Im just going to take a guess and say no

OpenStudy (lgbasallote):

right. because 2 - 2 is 0 and not 4 so (2,4) is NOT a solution (again, capitalizing for stress)

OpenStudy (anonymous):

oh well thank you so much :D